# Towards Reliable Forgetting: A Survey on Machine Unlearning Verification, Challenges, and Future Directions

### v1

@@ 2.2.3. Approximate Unlearning
Compared to exact unlearning, approximate unlearning relaxes the requirement of strict model equivalence. Specifically, it does not require the unlearned model $w^{\prime}$ to lie exactly within the hypothesis space $H_{r}=\{A(D_{r})\}$. Instead, the unlearning mechanism generates a model that approximates this target in one of the following ways:
@@
    Parameter-space approximation: The unlearned model $w^{\prime}$ lies within a hypothesis space $H^{\prime}$ that is close to $H_{r}$ in terms of model parameters.
@@
    Behavioral-space approximation: The unlearned model $w^{\prime}$ behaves similarly to models in $H_{r}$ on inputs related to the forgotten dataset $D_{f}$, thereby achieving approximate forgetting.
Parameter-space approximation. Parameter-space approximation aims to ensure that the unlearned model $w^{\prime}$ remains within a bounded distance from a model obtained via exact retraining. To achieve such bounded distance, researchers have proposed a range of methods, primarily including first-order approaches based on noise injection or projected gradients, and second-order approaches that utilize Hessian-based corrections.
@@
Second-order Methods. Second-order methods leverage the Hessian matrix or its approximations to more precisely adjust model parameters, aiming to approximate the retraining effect in parameter space. A representative example is the Newton Step correction mechanism proposed in (Guo et al., 2019), which updates the model weights using the inverse of the Hessian of the loss function, effectively canceling out the gradient contributions of the deleted data. Subsequent works (Sekhari et al., 2021; Jin et al., 2023) build upon this idea by developing Hessian estimation techniques that do not require access to the original training data, extending applicability to scenarios such as federated learning and recommender systems.
@@
Behavioral-space approximation. Behavioral-space approximation refers to ensuring that the unlearned model $w^{\prime}$ exhibits similar behavior to a retrained model when evaluated on specific inputs, even if their parameters differ. Specifically, these methods can be categorized into the following types.
@@
Fine-tuning (FT)-based methods. These methods construct carefully designed datasets and fine-tune the model so that it misclassifies the data intended to be forgotten, making the model behave as if it has never encountered those samples. Several works (Graves et al., 2021; Felps et al., 2020; Chen et al., 2023) propose modifying the labels of the data to be forgotten and fine-tuning the model on the resulting modified dataset. This process disrupts the association between the forgotten data and its correct label, allowing the model to gradually forget the data that no longer needs to be retained. Recently, to enhance privacy protection for forgotten data, some studies explore unlearning methods that do not require access to the forgotten data. Tarun et al. (2023) implement unlearning by fine-tuning using only a subset of the retained dataset and incorporating a specially constructed noise matrix. Chundawat et al. (2023b) extends the method from (Tarun et al., 2023) to apply to scenarios without the original data, further improving privacy protection.
@@
Gradient Ascent (GA)-based methods. These methods, as demonstrated by Wu et al. (2020); Liu et al. (2022a), apply gradient ascent to the loss computed on the data to be forgotten. Recently, (Kim et al., 2025) leverages the idea of gradient ascent to design an unlearning verification method tailored for text generation tasks in large multimodal models. This process degrades the model’s performance on that data, thereby driving its behavior closer to that of a model that has never encountered the forgotten data, and ultimately achieving approximate unlearning in the behavioral space.
@@
Influence-based methods. These methods aim to assess and eliminate the influence of specific training samples on model predictions or outputs. By leveraging influence functions to quantify how the presence of a data point affects the model’s output, they enable fine-grained control over the retention of training information. Golatkar et al. (2020a) propose a weight-sanitization approach that adjusts model parameters using the fisher information matrix (FIM). Golatkar et al. (2020b) introduce a neural tangent kernel linearization technique to more accurately estimate the direction of influence, and combine it with noise injection to mitigate residual effects. Building on influence function theory, Peste et al. (2021) present SSSE, a practical and scalable unlearning method. This method approximates the Hessian matrix using an empirical FIM and uses the Sherman-Morrison formula to efficiently compute its inverse, allowing the removal of the deleted sample’s influence without requiring access to the full training set. It is important to emphasize that these methods are fundamentally different from parameter-space approximation approaches, as they aim to eliminate the influence on model behavior rather than constrain differences in parameter space.
@@
Knowledge Distillation(KD)-based Methods. Knowledge distillation (Gou et al., 2021; Phuong and Lampert, 2019) is a technique that improves a student model by transferring behavioral patterns or soft targets from a more complex teacher model. In the context of machine unlearning, these methods treat the student as the unlearned model and achieve forgetting by blocking the transfer of behavior patterns from the teacher that are associated with the data to be forgotten. Chundawat et al. (2023b) introduce a band-pass filter to remove data containing forgotten information, ensuring that the student model only receives the retained information. Chundawat et al. (2023a) establish a teacher-student framework with both competent and incompetent teachers, using selective knowledge transfer to enable the student model to forget certain information. Kurmanji et al. (2023) present the SCRUB method, which effectively achieves forgetting by minimizing KL divergence and task loss on non-forgotten datasets while maximizing KL divergence on forgotten datasets.
@@
Membership Game-based Methods. These methods focus on aligning membership-related behavior, aiming to ensure that the final model exhibits membership inference responses on the forgotten dataset $D_{f}$ that are indistinguishable from those of a model that has never encountered $D_{f}$. In doing so, they seek to eliminate observable behavioral traces of the forgotten data. Recent works (Tu et al., 2024; Di et al., 2024; Sharma et al., 2024) formalize this process as a game between the unlearning mechanism and an adversary, where the model’s behavior is optimized to evade membership inference attacks. However, these methods rely exclusively on behavioral signals related to membership privacy over the forgotten dataset, and their effectiveness heavily depends on the strength of the specific attack strategy used. As a result, their reliability and generalizability are highly sensitive to the choice and quality of the underlying inference mechanism.
@@
Methods Scratch SISA Non-SISA First-order Second-order FT, GA Influence, KD Membership games
@@ 2.3. Machine Unlearning Verification
Figure 3. The process of machine unlearning verification
@@
$V(w^{\prime},D_{f})\in\{0,1\}$
where $V(w^{\prime},D_{f})=1$ indicates that unlearning has succeeded, meaning no residual influence of $D_{f}$ is detectable in the model $w^{\prime}$; whereas $V(w^{\prime},D_{f})=0$ denotes that the unlearning process has failed to fully remove the impact of $D_{f}$.
@@
Who participates in the verification process? The unlearning verification process typically involves three main entities: the data owner, who requests the removal of a specified subset $D_{f}$ from the training data; the model provider, who applies an unlearning mechanism and returns a model $w^{\prime}$ that is claimed to have forgotten $D_{f}$; and the auditor, who evaluates whether $w^{\prime}$ still retains information about the forgotten data, as illustrated in Figure 3.
@@
Verification, fill=purple!20 [Fingerprint Verification, fill=purple!10 [Passive Fingerprint.
@@
Verification, fill=purple!10 [Data Inference(Tu et al., 2024; Di et al., 2024; Kim et al., 2025)
@@
(Golatkar et al., 2020b, a; Peste et al., 2021; Jeon et al., 2024b; Vidal et al., 2024) ] [Model Sensitivity (Zhou et al., 2025a; Becker and Liebig, 2022)] ] ] [Parametric Unlearning
@@
(Guo et al., 2019; Sekhari et al., 2021; Jin et al., 2023; Zhang et al., 2024b; Qiao et al., 2024)
(Liu et al., 2023a; Chien et al., 2022; Dong et al., 2024; Wu et al., 2023; Yi and Wei, 2024)] ] [Reproducible Verification, fill=orange!10 [PoUL(Thudi et al., 2022; Zhang et al., 2024a)] [PoUL with Cryptography
(Weng et al., 2024; Eisenhofer et al., 2022)] ] ] ]
@@
  • Behavioral unlearning verification. Behavioral unlearning verification aims to determine whether an unlearned model has successfully forgotten the target data by evaluating its behavior on specific inputs. These methods define a model’s behavioral signature and assess the predicate $V(w^{\prime},D_{f})$ based on observable behavioral changes. In particular, the evaluation focuses on the model’s behavior $\mathcal{B}(w^{\prime},{Q}_{D_{f}}(D_{f}))$ over a designated query set ${Q}_{D_{f}}(D_{f})$, which consists of inputs related to the forgotten data $D_{f}$. To verify unlearning, the auditor compares this behavior to a reference behavior set ${T}_{\neg D_{f}}$, representing the expected behavior of models that have never seen $D_{f}$. If $\mathcal{B}(w^{\prime},{Q}_{D_{f}}(D_{f}))$ aligns closely with ${T}_{\neg D_{f}}$, the model is considered to have forgotten the data.
  • Parametric unlearning verification. These methods evaluate whether unlearning has been achieved by measuring the difference in parameter space between the target model and a reference model that has never been exposed to the forgotten data. If the difference falls below a predefined threshold, the model is considered to meet the unlearning criterion. These approaches rely on specific theoretical assumptions, as discussed in Section 4. Compared to behavioral verification, parameter-based verification typically offers greater reliability. However, it often requires access to internal model parameters or training dynamics, and may incur higher computational costs.These methods validate forgetting by assessing the difference between the parameters of the target unlearned model and the parameter space $H^{\prime}$ of models that have never seen the forgotten data. Representative approaches include differential privacy verification and reproducibility verification. Compared to behavioral methods, parametric verification typically provides more reliable results. However, it generally requires access to internal model parameters or training dynamics, and often incurs higher computational costs.
@@
3. Behavioral Unlearning Verification.
Existing behavioral verification methods can be classified into three categories: fingerprint methods, privacy analysis methods, and model performance methods. Their behavioral constructions are summarized in Table 4.
@@
Intuition Models not trained on $D_{f}$ do not show its fingerprints Models not trained on $D_{f}$ should not leak its privacy Models trained with and without $D_{f}$ perform differently on $D_{f}$.
Types of Behavior Passive Fingerprint, Active Fingerprint Data Inference, Reconstruction, Weight analysis Accuracy, Relearn Time, Model Output, Model Gradient
@@ 3.1. Fingerprint
In machine learning, fingerprints refer to a specific response pattern that a model develops as a result of having been exposed to certain training data. This pattern typically manifests in the model’s behavior when processing particular inputs.
Fingerprint verification methods construct a query set ${Q}_{D_{f}}$ to reveal whether the model still retains fingerprint information associated with the forgotten data $D_{f}$. The model’s responses to these queries are denoted as $\mathcal{B}$, indicating whether any residual fingerprints are exposed during evaluation. The rationale behind the behavior set ${T}_{\neg D_{f}}$ is grounded in a simple intuition: a model that has never encountered $D_{f}$ should not exhibit distinctive responses (fingerprints) to inputs related to it. Therefore, if a model that claims to have unlearned $D_{f}$ still exhibits fingerprint-like behavior on the query set, it suggests that the model has not passed the unlearning verification.
Fingerprint verification methods can be categorized as active or passive, based on whether the unlearned data has undergone prior proactive intervention.
@@
Adversarial example techniques refer to methods for generating inputs with imperceptible perturbations that intentionally trigger incorrect or unstable outputs in machine learning models (Goodfellow et al., 2014; Brendel et al., 2017; Cheng et al., 2019; Zhou et al., 2024a, 2023b, 2023a; Li et al., 2024b; Zhou et al., 2025b, 2024b; Wang et al., 2025a; Song et al., 2025). Although originally developed to test model robustness and to design attacks, these techniques have increasingly been repurposed for auxiliary objectives such as unlearning verification. By using adversarially perturbed inputs as diagnostic probes, researchers can assess whether a model remains sensitive to specific data points that are supposed to have been erased.
Xuan and Li (2025) propose a white-box adversarial verification method for machine unlearning, termed Unlearning Mapping Attack (UMA), which is applicable to both classification and generative models. In classification tasks, UMA perturbs known forget samples to check whether the model still produces the same predictions as before unlearning. If it does, this suggests that the model retains latent representations associated with the forgotten class, indicating incomplete unlearning. This residual sensitivity implies that decision boundaries related to the forgotten data have not been effectively removed. In generative tasks, UMA applies adversarial perturbations to masked inputs to induce the reconstruction of forgotten content. If the model is still capable of producing outputs that resemble the original data, it demonstrates that structural knowledge of the forgotten samples persists, further undermining the effectiveness of the unlearning process.
Similarly, Zhang et al. (2024d) propose UnlearnDiffAtk, which constructs adversarial prompts specifically for diffusion models. If the model, after unlearning, still generates forgotten content in response to these prompts, it is considered to have failed the unlearning verification.
@@
Active fingerprint verification involves constructing or modifying specific data samples related to the unlearning target before or during training, in order to embed identifiable fingerprint information. Several active fingerprint methods construct verifiable fingerprints for unlearning verification by strategically modifying the target data to be forgotten, such that only models exposed to the data can correctly identify them. For instance, Goel et al. (Goel et al., 2022) propose Inter-class Confusion Testing, which generates adversarially perturbed samples during training and later evaluates the model’s ability to recognize these samples after unlearning. If the model fails to recognize them, the unlearning is considered successful. In the context of federated learning, Tam et al. (Tam et al., 2024) introduce adversarial noise as validation markers to assess whether information from a specific domain has been effectively removed. Similarly, Gao et al. (Gao et al., 2024) explore the use of label flipping and adversarial perturbations to construct verification signals for unlearning.
@@
In addition, several studies have demonstrated the feasibility of using backdoor mechanisms for unlearning verification. Essentially, a backdoor attack is a form of data poisoning in which the attacker injects trigger-embedded samples into the training data (Guo et al., 2023; Hu et al., 2022; Wang et al., 2024b; Zhang et al., 2024c, 2025). Once compromised, the model behaves normally on clean data but produces targeted incorrect predictions when presented with inputs containing the trigger. This attack paradigm can be repurposed for unlearning verification: the model owner deliberately implants backdoor triggers into the model, and after executing unlearning algorithms, checks whether the model’s predictions on the trigger inputs have changed. If they remain the same, unlearning is deemed to have failed; otherwise, it is considered successful. Recently, studies (Guo et al., 2023; Han et al., 2025) have proposed backdoor-based verification strategies, in which specific triggers are embedded into target data during training. In the verification phase, the model’s response to these trigger-embedded queries is monitored to determine whether it still predicts them as the user-specified target class. This approach enables black-box unlearning verification without requiring access to the model’s internal parameters. Zhang et al. (2024e) propose DuplexGuard, which associates each forget subset with a corresponding emergent subset. The model exhibits no backdoor behavior before unlearning, but once the forget subset is removed, the backdoor effect on the emergent subset becomes active, serving as a verification signal. In study (Sommer et al., 2022), users inject unique backdoor triggers into their own data, and the model’s response to these triggers after deletion requests is statistically evaluated using hypothesis testing. This provides a quantitative and high-confidence verification mechanism that remains robust against state-of-the-art backdoor defenses and adaptive server behaviors. Xu et al. [108] propose IndirectVerify, a novel unlearning verification scheme that constructs influential sample pairs, consisting of a trigger sample and a corresponding reaction sample. By perturbing the trigger sample to influence the classification of the reaction sample via gradient matching, the method ensures that successful unlearning modifies the prediction of the reaction sample.
@@
In addition to backdoor,Izzo et al. (2021) propose the Feature Injection Test to evaluate unlearning effectiveness. The method introduces an artificial feature into the target data that is activated for only a small subset of samples and is highly predictive of the label. During training, the model assigns a large weight to this feature due to its strong correlation with the output. If the target samples are successfully deleted, an effective unlearning mechanism should cause the weight associated with this feature to drop significantly. Lu et al. (Lu et al., 2025) propose WaterDrum, a black-box unlearning verification method for large language models that adopts an active fingerprinting paradigm. It embeds robust text watermarks, parameterized by private cryptographic keys, into the training data. Models trained on such data emit outputs containing identifiable signals, which vanish once the data is successfully unlearned. Verification is conducted privately by the data owner using their key to detect the presence or absence of the watermark.

### v2

@@ 2.2.3. Approximate Unlearning
Compared to exact unlearning, approximate unlearning relaxes the requirement of strict model equivalence. Specifically, it does not require the unlearned model $\mathbf{w}^{\prime}$ to lie exactly within the hypothesis space $H_{r}=\{A(D_{r})\}$. Instead, the unlearning mechanism generates a model that approximates this target in one of the following ways:
@@
    Parameter-space approximation: The unlearned model $\mathbf{w}^{\prime}$ lies within a hypothesis space $H^{\prime}$ that is close to $H_{r}$ in terms of model parameters.
@@
    Behavioral-space approximation: The unlearned model $\mathbf{w}^{\prime}$ behaves similarly to models in $H_{r}$ on inputs related to the forgotten dataset $D_{f}$, thereby achieving approximate forgetting.
Parameter-space approximation. Parameter-space approximation aims to ensure that the unlearned model $\mathbf{w}^{\prime}$ remains within a bounded distance from a model obtained via exact retraining. To achieve such bounded distance, researchers have proposed a range of methods, primarily including first-order approaches based on noise injection or projected gradients, and second-order approaches that utilize Hessian-based corrections.
@@
Second-order Methods. Second-order methods leverage the Hessian matrix or its approximations to more precisely adjust model parameters, aiming to approximate the retraining effect in parameter space. A representative example is the Newton Step correction mechanism proposed in (Guo et al., 2020), which updates the model weights using the inverse of the Hessian of the loss function, effectively canceling out the gradient contributions of the deleted data. Subsequent works (Sekhari et al., 2021; Jin et al., 2023) build upon this idea by developing Hessian estimation techniques that do not require access to the original training data, extending applicability to scenarios such as federated learning and recommender systems.
@@
Behavioral-space approximation. Behavioral-space approximation refers to ensuring that the unlearned model $\mathbf{w}^{\prime}$ exhibits similar behavior to a retrained model when evaluated on specific inputs, even if their parameters differ. Specifically, these methods can be categorized into the following types.
@@
Fine-tuning (FT)-based Methods. These methods construct carefully designed datasets and fine-tune the model so that it misclassifies the data intended to be forgotten, making the model behave as if it has never encountered those samples. Several works (Graves et al., 2021; Felps et al., 2020; Chen et al., 2023) proposed modifying the labels of the data to be forgotten and fine-tuning the model on the resulting modified dataset. This process disrupts the association between the forgotten data and its correct label, allowing the model to gradually forget the data that no longer needs to be retained. Recently, to enhance privacy protection for forgotten data, some studies explore unlearning methods that do not require access to the forgotten data. Tarun et al. (2023) implemented unlearning by fine-tuning using only a subset of the retained dataset and incorporating a specially constructed noise matrix. Chundawat et al. (2023b) extended the method from (Tarun et al., 2023) to apply to scenarios without the original data, further improving privacy protection.
@@
Gradient Ascent (GA)-based Methods. These methods, as demonstrated by (Wu et al., 2020; Liu et al., 2022a), apply gradient ascent to the loss computed on the data to be forgotten. Recently, Kim et al. (2025) leverages the idea of gradient ascent to design an unlearning verification method tailored for text generation tasks in large multimodal models. This process degrades the model’s performance on that data, thereby driving its behavior closer to that of a model that has never encountered the forgotten data, and ultimately achieving approximate unlearning in the behavioral space.
@@
Influence-based Methods. These methods aim to assess and eliminate the influence of specific training samples on model predictions or outputs. By leveraging influence functions or the Fisher Information Matrix (FIM) to quantify the impact of individual data points on models, these methods enable fine-grained control over the retention of training information. Golatkar et al. (2020a) proposed a weight-sanitization approach that adjusts model parameters using the fisher information matrix. Golatkar et al. (2020b) introduced a neural tangent kernel linearization technique to more accurately estimate the direction of influence, and combine it with noise injection to mitigate residual effects. Building on influence function theory, Peste et al. (2021) presented SSSE, a practical and scalable unlearning method. This method approximates the Hessian matrix using an empirical FIM and uses the Sherman-Morrison formula to efficiently compute its inverse, allowing the removal of the deleted sample’s influence without requiring access to the full training set. It is important to emphasize that these methods are fundamentally different from parameter-space approximation approaches, as they aim to eliminate the influence on model behavior rather than constrain differences in parameter space.
@@
Knowledge Distillation (KD)-based Methods. Knowledge distillation (Gou et al., 2021; Phuong and Lampert, 2019) is a technique that improves a student model by transferring behavioral patterns or soft targets from a more complex teacher model. In the context of machine unlearning, these methods treat the student as the unlearned model and achieve forgetting by blocking the transfer of behavior patterns from the teacher that are associated with the data to be forgotten. Chundawat et al. (2023b) introduced a band-pass filter to remove data containing forgotten information, ensuring that the student model only receives the retained information. Chundawat et al. (2023a) established a teacher-student framework with both competent and incompetent teachers, using selective knowledge transfer to enable the student model to forget certain information. Kurmanji et al. (2023) presented the SCRUB method, which effectively achieves forgetting by minimizing KL divergence and task loss on non-forgotten datasets while maximizing KL divergence on forgotten datasets.
@@
Membership Game-based Methods. These methods focus on aligning membership-related behavior, aiming to ensure that the final model exhibits membership inference responses on the forgotten dataset $D_{f}$ that are indistinguishable from those of a model that has never encountered $D_{f}$. In doing so, they seek to eliminate observable behavioral traces of the forgotten data. Recent works (Tu et al., 2024; Di et al., 2025; Sharma et al., 2024) formalized this process as a game between the unlearning mechanism and an adversary, where the model’s behavior is optimized to evade membership inference attacks. However, these methods rely exclusively on behavioral signals related to membership privacy over the forgotten dataset, and their effectiveness heavily depends on the strength of the specific attack strategy used. As a result, their reliability and generalizability are highly sensitive to the choice and quality of the underlying inference mechanism.
@@
Methods Scratch, SISA, Non-SISA First-order, Second-order FT, GA, Influence, KD, Membership games
@@ 2.3. Machine Unlearning Verification
Figure 3. The process of machine unlearning verification.
@@
$V(\mathbf{w}^{\prime},D_{f})\in\{0,1\},$
where $V(\mathbf{w}^{\prime},D_{f})=1$ indicates that unlearning has succeeded, meaning no residual influence of $D_{f}$ is detectable in the model $\mathbf{w}^{\prime}$; whereas $V(\mathbf{w}^{\prime},D_{f})=0$ denotes that the unlearning process has failed to fully remove the impact of $D_{f}$.
@@
Who participates in the verification process? The unlearning verification process typically involves three main entities: the data owner, who requests the removal of a specified subset $D_{f}$ from the training data; the model provider, who applies an unlearning mechanism and returns a model $\mathbf{w}^{\prime}$ that is claimed to have forgotten $D_{f}$; and the auditor, who evaluates whether $\mathbf{w}^{\prime}$ still retains information about the forgotten data, as illustrated in Figure 3.
@@
Verification, fill=purple!20 [Fingerprint Verification, fill=purple!10 [Passive Fingerprint
@@
Verification, fill=purple!10 [Data Inference (Tu et al., 2024; Di et al., 2025; Kim et al., 2025)
@@
(Golatkar et al., 2020b, a; Chundawat et al., 2023a; Nguyen et al., 2020; Peste et al., 2021; Jeon et al., 2024; Vidal et al., 2024) ] [Model Sensitivity (Zhou et al., 2025a; Becker and Liebig, 2022)] ] ] [Parametric Unlearning
@@
(Guo et al., 2020; Sekhari et al., 2021; Jin et al., 2023; Zhang et al., 2024b; Qiao et al., 2024)
(Liu et al., 2023a; Chien et al., 2022; Dong et al., 2024; Wu et al., 2023; Yi and Wei, 2024)] ] [Reproducible Verification, fill=orange!10 [PoUL (Thudi et al., 2022; Zhang et al., 2024a; Wu et al., 2020; Nguyen et al., 2022b)] [PoUL with Cryptography
(Weng et al., 2024; Eisenhofer et al., 2025)] ] ] ]
@@
  1. (1)
    Behavioral unlearning verification. Behavioral unlearning verification aims to determine whether a model has successfully forgotten the $D_{f}$ by evaluating its responses to specific inputs. Such methods typically define behavioral characteristics of the model and assess a predicate $V(\mathbf{w}^{\prime},D_{f})$ based on observable changes in these behaviors. Specifically, the evaluation focuses on the model’s behavior $\mathcal{B}(\mathbf{w}^{\prime},Q_{D_{f}})$ over a designated query set $Q_{D_{f}}$, which contains inputs related to the forgotten data $D_{f}$. If this behavior is consistent with or sufficiently close to that of $\mathcal{B}(\mathbf{w}^{*},Q_{D_{f}})$, where $\mathbf{w}^{*}$ is a model that has never seen $D_{f}$, the unlearned model is considered to have successfully forgotten $D_{f}$.
  2. (2)
    Parametric unlearning verification. These methods evaluate whether unlearning has been achieved by measuring the difference in parameter space between the target model and a reference model that has never been exposed to the forgotten data. If the difference falls below a predefined threshold, the model is considered to meet the unlearning criterion. These approaches rely on specific formal assumptions, as discussed in Section 4. Compared to behavioral verification, parametric verification typically offers greater reliability. However, it often requires access to internal model parameters or training dynamics, and may incur higher computational costs.
@@
3. Behavioral Unlearning Verification
Existing behavioral verification methods can be classified into three categories: fingerprint methods, privacy analysis methods, and model performance methods. These approaches differ in the specific types of model behaviors they exploit as verification signals, as well as in the intuitions underlying those behaviors. Table 4 summarizes these behavioral foundations, outlining the distinct principles and observations upon which each method relies.
@@ 3. Behavioral Unlearning Verification
Intuition Models not trained on $D_{f}$ do not show its fingerprints Models not trained on $D_{f}$ should not leak its privacy Models trained with and without $D_{f}$ perform differently on $D_{f}$
Types of Behavior Passive Fingerprint, Active Fingerprint Data Inference, Data Reconstruction Task Quality, Relearn Time, Model Output, Model Sensitivity
@@ 3.1. Fingerprint
In machine learning, fingerprints refer to distinctive response patterns a model acquires from exposure to specific training data (Dathathri et al., 2018; Chen et al., 2018). These patterns can be detected through a query set $Q_{D_{f}}$, which probes whether the model retains behavioral traces associated with the forgotten data $D_{f}$. The model’s responses $\mathcal{B}$ reveal whether such residual fingerprints remain.
The core intuition is that models trained on $D_{f}$ often exhibit systematic deviations from those never exposed to it, for example, consistently misclassifying triggered inputs into a specific class or responding differently to adversarially perturbed inputs. Fingerprint verification audits such signals to assess whether $D_{f}$ has been truly forgotten. Fingerprint verification methods can be categorized as active or passive, based on whether the unlearned data has undergone prior proactive intervention.
@@
Adversarial example techniques refer to methods for generating inputs with imperceptible perturbations that intentionally trigger incorrect or unstable outputs in machine learning models (Zhou et al., 2024a, 2023b, 2023a; Li et al., 2024b; Zhou et al., 2025b, 2024b; Wang et al., 2025a; Song et al., 2025). Although originally developed to test model robustness and to design attacks, these techniques have increasingly been repurposed for auxiliary objectives such as unlearning verification. By using adversarially perturbed inputs as diagnostic probes, researchers can assess whether a model remains sensitive to specific data points that are supposed to have been erased.
Xuan and Li (2025) proposed a white-box adversarial verification method for machine unlearning, termed Unlearning Mapping Attack (UMA), which is applicable to both classification and generative models. In classification tasks, UMA perturbs known forget samples to check whether the model still produces the same predictions as before unlearning. If it does, this suggests that the model retains latent representations associated with the forgotten class, indicating incomplete unlearning. This residual sensitivity implies that decision boundaries related to the forgotten data have not been effectively removed. In generative tasks, UMA crafts adversarial perturbations on the masked inputs (i.e., partially occluded versions) of samples from the forget set. The goal is to induce the unlearned generative model to produce outputs that closely resemble those generated by the original model. Successful reconstruction indicates that the model still retains residual information about the forgotten data.
Similarly, Zhang et al. (2024d) proposed UnlearnDiffAtk, which constructs adversarial prompts specifically for diffusion models. If the unlearned model still generates forgotten content in response to these prompts, it is considered to have failed the unlearning verification.
@@
Active fingerprint verification involves constructing or modifying specific data samples related to the unlearning target before or during training, in order to embed identifiable fingerprint information. Several active fingerprint methods construct verifiable fingerprints for unlearning verification by strategically modifying the target data to be forgotten, such that only models exposed to the data can correctly identify them. For instance, Goel et al. (Goel et al., 2022) proposed Inter-class Confusion Testing, which generates adversarially perturbed samples during training and later evaluates the model’s ability to recognize these samples after unlearning. If the model fails to recognize them, the unlearning is considered successful. In the context of federated learning, Tam et al. (Tam et al., 2024) introduced adversarial noise as validation markers to assess whether information from a specific domain has been effectively removed. Similarly, Gao et al. (Gao et al., 2024) explored the use of label-flipping techniques to construct verification signals for unlearning.
@@
In addition, several studies have demonstrated the feasibility of using backdoor mechanisms for unlearning verification. Essentially, a backdoor attack is a form of data poisoning in which the attacker injects trigger-embedded samples into the training data (Guo et al., 2023; Hu et al., 2022; Wang et al., 2024b; Zhang et al., 2024c, 2025). Once backdoored, the model behaves normally on clean data but produces targeted incorrect predictions when presented with inputs containing the trigger. This attack paradigm can be repurposed for unlearning verification: the model owner deliberately implants backdoor triggers into the model, and after executing unlearning algorithms, checks whether the model’s predictions on the trigger inputs have changed. If they remain the same, unlearning is deemed to have failed; otherwise, it is considered successful. Recently, studies (Guo et al., 2023; Han et al., 2025) have proposed backdoor-based verification strategies, in which specific triggers are embedded into target data during training. In the verification phase, the model’s response to these trigger-embedded queries is monitored to determine whether it still predicts them as the user-specified target class. This approach enables black-box unlearning verification without requiring access to the model’s internal parameters. Zhang et al. (2024e) proposed DuplexGuard, which associates each forget subset with a corresponding emergent subset. The model exhibits no backdoor behavior before unlearning, but once the forget subset is removed, the backdoor effect on the emergent subset becomes active, serving as a verification signal. In study (Sommer et al., 2022), users inject unique backdoor triggers into their own data, and the model’s response to these triggers after deletion requests is statistically evaluated using hypothesis testing. This provides a quantitative and high-confidence verification mechanism that remains robust against state-of-the-art backdoor defenses and adaptive server behaviors. Xu et al. (2024b) proposed IndirectVerify, a novel unlearning verification scheme that constructs influential sample pairs, consisting of a trigger sample and a corresponding reaction sample. By perturbing the trigger sample to influence the classification of the reaction sample via gradient matching, the method ensures that successful unlearning modifies the prediction of the reaction sample.
@@
In addition to backdoor-based approaches, Izzo et al. (2021) proposed a feature injection test to evaluate the effectiveness of unlearning. This method introduces an artificial feature into $D_{f}$ that is activated only for a small subset of samples and is highly predictive of the label. During training, the model typically develops a strong response to this feature. Therefore, if the unlearning mechanism is effective, removing the associated data should lead to a significant reduction in the model’s reliance on this feature, which can manifest as changes in predictive behavior. Lu et al. (2025) proposed WaterDrum, a black-box unlearning verification method for large language models that adopts an active fingerprinting paradigm. It embeds robust text watermarks, parameterized by private cryptographic keys, into the training data. Models trained on such data emit outputs containing identifiable signals, which vanish once the data is successfully unlearned. Verification is conducted privately by the data owner using their key to detect the presence or absence of the watermark.